OpenStudy (decarr432):

let f(x)=1/x and g(x)=x^2-9x What two numbers are not in the domain of (f o g)

OpenStudy (astrophysics):

Lets start off by setting it up as f(g(x)) first, go ahead and do that

OpenStudy (decarr432):

1/(x^2-9x) right?

OpenStudy (astrophysics):

Perfect, \[f(g(x)) = \frac{ 1 }{ x^2-9x } = \frac{ 1 }{ x(x-9) }\] I simplified it a bit more for you so it's easier to see (by factoring out an x).

OpenStudy (astrophysics):

Good so far?

OpenStudy (decarr432):

Yeah

OpenStudy (astrophysics):

Ok now we can see the restrictions by \[x^2-9x \neq 0 \] or you can see it clearly in the simplified form by setting \[x \neq 0 ~~~\text{and}~~~x -9 \neq 0\] I sort of gave away one of the numbers not allowed haha but it's ok long as you understand it.

OpenStudy (astrophysics):

Again we solve for x

OpenStudy (astrophysics):

The reason I have set it cannot equal to 0 is because if we divide over 0 we get undefined right?

OpenStudy (decarr432):

Yeah makes sense

OpenStudy (decarr432):

greater than zero right?

OpenStudy (decarr432):

Could it be a negative number?

OpenStudy (astrophysics):

Not quite, we're not working with square roots, lets think about this. So we have one of our restrictions as I showed you which was \[x \neq 0\] now lets see what happens if we plug in the 0 in our function. \[f(g(0)) = \frac{ 1 }{ 0(x-9) } = \frac{ 1 }{ 0 }\] uh oh undefined..that's not going to work, so it will be one of our two numbers not in the domain.

OpenStudy (astrophysics):

For the second number \[x-9 \neq 0\] solve for x

OpenStudy (decarr432):

x can be any number as long as the outcome does not equal zero correct?

OpenStudy (astrophysics):

Right

OpenStudy (decarr432):

Okay so how do we figure out two separate numbers that are not in the domain?

OpenStudy (decarr432):

one would be 9 right?

OpenStudy (astrophysics):

Right and the other is 0 we got that from (quoting myself lol) \(\color{blue}{\text{Originally Posted by}}\) @Astrophysics Ok now we can see the restrictions by \[x^2-9x \neq 0 \] or you can see it clearly in the simplified form by setting \[x \neq 0 ~~~\text{and}~~~x -9 \neq 0\] I sort of gave away one of the numbers not allowed haha but it's ok long as you understand it. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (astrophysics):

We looked at our function and saw our denominator and set that equal to 0

OpenStudy (decarr432):

Okay so recap

OpenStudy (decarr432):

Any number is allowed as long as it doesn't come out to be zero in the end?

OpenStudy (astrophysics):

0 in the `denominator`

OpenStudy (astrophysics):

because we get undefined

OpenStudy (decarr432):

yeah but im scared to submit my answer lol

OpenStudy (astrophysics):

Why? We know our two numbers that are not in the domain, I suggest you go over everything in this post again

OpenStudy (decarr432):

because it says g(x) x^2-9x

OpenStudy (astrophysics):

That's ok

OpenStudy (decarr432):

its the square thats confusing me

OpenStudy (decarr432):

Well I manned up and it was correct thanks bud again lol I owe you alot haha

OpenStudy (astrophysics):

No worries, just go over it again, don't let math scare you ;)! They will try to trick you but it's the same concept being applied over and over again

OpenStudy (decarr432):

Okay last one I'll submit the question